OpenStudy (anonymous):

Solve the equation - 4 = (square root of p) – 2

OpenStudy (anonymous):

\[4 = \sqrt{p} - 2\]

OpenStudy (anonymous):

p=36

OpenStudy (anonymous):

thanks :) how did you get that?

OpenStudy (anonymous):

√p = 6, because 6-2 = 4....:) so p= 6^2

OpenStudy (anonymous):

ohh ok got it... thanks again